# breaklines Repository-Level Compositional Code Translation and Validation

### v2

@@ 7. Evaluation
Translation Bugs and Fixes. How much effort do developers spend completing the partial translations created by AlphaTrans and achieving all passing tests? What is the nature of translation bugs?
@@
Impact of Test Decomposition. To what extent does test decomposition impact the validation results?
@@ 7.1. Experiment Setup
We followed three steps for selecting subjects:
1- Mining. We mined GitHub and retrieved a list of repositories that use Java as the primary language, are self-contained (include build files, etc.), and have more than stars with at least one commit pushed within the last months.
2- Filtering. We filtered out projects based on the number of call edges in their call graphs: removed those with less than call edges to ensure the subject projects are big enough to challenge AlphaTrans. We also removed those with more than call edges to reduce the computation and manual effort for further steps. Per GraalVM requirements, we only selected projects we could successfully build (compile and achieve green tests) using Java at .
3- Reduction. AlphaTrans currently supports the following Java APIs: core Java API (java.util, java.text, java.lang, java.io, java.nio, java.net, java.time, and java.math) and third-party libraries (org.opentest4j, org.slf4j.Logger, and org.junit). We automatically removed all other third-party library dependencies and their usage in the source code in the selected projects. We chose a project if at least of its total methods were preserved after such process. Table 1 shows the list of ten projects used in the evaluation of AlphaTrans and details about their size (classes, methods, tests, and fragments).
AlphaTrans uses CodeQL (GitHub, 2024) and tree-sitter (Tree-Sitter, 2024) for static analysis. For running tests, validating translation, and computing coverage, AlphaTrans uses GraalVM (Oracle, 2024), JUnit 4 and 5 (Team, 2024b), Pytest 8.2.1 (pytest dev, 2024), JaCoCo (Team, 2024a), and Python’s coverage (Batchelder, 2024). AlphaTrans can work with API-access and open-access LLMs. We considered the following criteria for selecting the LLM: (1) for better reproducibility, we prioritized open-access models; (2) due to computing constraints, we wanted an LLM with moderate size ( but parameters); (3) the model should perform reasonably well in code-related tasks; and (4) the model should have fast inference time due to the huge number of prompts. Per the mentioned criteria, we selected DeepSeek-Coder-33b-Instruct for our experiments. We prompted DeepSeek-Coder-33b-Instruct under temperature setting to make translations reproducible. We used the LLM’s default setting for other parameters. For the base prompting (Algorithm 2), we set the minimum and maximum values of the reprompting budget to and . For the feedback prompting, we set the reprompting budget to , i.e., AlphaTrans attempts to fix issues with feedback for only once.
@@ 7.2.1. Effectiveness in type resolution and skeleton validation:
| Subjects | # Classes | # Methods | # Original Tests | Method Coverage (%) | ATR (%) | SV (%) | # Fragments | |||
| Fields | Application Methods | Test Methods | ||||||||
| Application | Test | |||||||||
| cli (Foundation, 2024a) | 58 | 664 | 437 | 93.4 | 96.6 | 100 | 104 | 57 | 273 | 2,180 |
| codec (Foundation, 2024b) | 156 | 1780 | 992 | 86.5 | 96.0 | 100 | 425 | 140 | 680 | 2,849 |
| csv (Foundation, 2024c) | 41 | 694 | 309 | 87.7 | 92.3 | 100 | 146 | 35 | 235 | 1,272 |
| exec (Foundation, 2024d) | 56 | 407 | 70 | 53.2 | 78.9 | 100 | 104 | 27 | 248 | 327 |
| fast-pfor (Lemire, 2024) | 82 | 971 | 82 | 50.5 | 87.4 | 100 | 127 | 14 | 748 | 302 |
| fileupload (Foundation, 2024e) | 49 | 381 | 39 | 20.3 | 98.3 | 100 | 121 | 39 | 192 | 194 |
| graph (Foundation, 2024f) | 118 | 879 | 146 | 58.4 | 97.1 | 100 | 216 | 29 | 541 | 977 |
| jansi (Source, 2024) | 48 | 474 | 107 | 25.4 | 84.8 | 100 | 378 | 0 | 409 | 123 |
| pool (Foundation, 2024g) | 98 | 1097 | 73 | 24.6 | 91.6 | 100 | 203 | 91 | 682 | 649 |
| validator (Foundation, 2024h) | 130 | 1,228 | 464 | 61.9 | 95.5 | 100 | 421 | 209 | 646 | 1,463 |
| Total/Average | 836 | 8,575 | 2,719 | 56.2 | 91.9 | 100 | 2,245 | 641 | 4,654 | 10,336 |
@@ 7.2.2. Effectiveness in compositional translation and validation:
Table 2 shows the detailed compositional translation and validation results. The AMF column indicates the total number of application method fragments. The numbers in subsequent columns demonstrate the effectiveness of AlphaTrans in translation and validation of AMFs only444Due to space limit, please refer to our artifact (Lab, 2024) for details about all translation and validation of other fragments shown in Table 1.. The Syntax Check column indicates the percentage of AMFs that pass syntactic validation. Column SNEF shows the percentage of AMFs not covered by source PL tests. These results show that AlphaTrans successfully creates syntactically correct code (, on average). We also observe that of AMFs are not covered during the execution of any test, i.e., we cannot go beyond syntactic check and validate their runtime behavior or functional equivalence.
@@
For of AMFs that can be covered by source project tests, AlphaTrans attempts to validate the functional equivalence using GraalVM. The super column GraalVM shows the percentage of AMFs that GraalVM executes and successfully validates (GS), executes but there is a test assertion failure (GF), and cannot execute due to its limitation (GE) mentioned in §6.1. On average, (min= and max=), (min= and max=), and (min= and max=) of AMFs resulted in Graal Success, Graal Fail, and Graal Error, respectively. Note that these numbers add up to of AMFs that were covered by source program tests. With respect to only covered methods, GraalVM Success will be . Furthermore, our analysis shows that a high portion of methods that are not covered by tests are either abstract methods or getter/setter methods. If their translations are syntactically correct, they are also likely functionally equivalent, which can ramp up the success rate (§7.5). Spearman Rank Order Correlation (Spearman, 1961) indicates a strong positive correlation between the method coverage (Table 1) and GS numbers (), confirming that with better method coverage, the validated AMFs are very likely to be higher.
| Subjects | AMF | Syntax Check (%) | SNEF (%) | GraalVM | Test Translation | M1 | ||||||||||||||||
| GS (%) | GF (%) | GE (%) | TNEF (%) | ATP (%) | OTF (%) | MTF (%) | ATF (%) | TPR (%) | ||||||||||||||
| Overall | RE | AF | Overall | RE | AF | Overall | RE | AF | All | Some | ||||||||||||
| cli | 273 | 100 | 6.6 | 70.7 | 11.7 | 11.0 | 25.6 | 10.3 | 12.8 | 51.4 | 48.6 | 35.9 | 91.0 | 9.0 | 8.8 | 100 | 0.0 | 10.1 | 0 | 14 | ||
| codec | 680 | 98.5 | 13.5 | 25.4 | 49.9 | 11.2 | 61.0 | 4.1 | 4.4 | 60.0 | 40.0 | 12.8 | 55.1 | 44.9 | 4.1 | 75.2 | 24.8 | 9.4 | 11 | 24 | ||
| csv | 235 | 98.7 | 12.3 | 38.7 | 26.8 | 22.1 | 83.0 | 0.9 | 1.3 | 33.3 | 66.7 | 0.0 | 0.0 | 0.0 | 2.6 | 100 | 0.0 | 0.2 | 0 | 0 | ||
| exec | 248 | 100 | 46.8 | 33.5 | 2.0 | 17.7 | 31.5 | 4.4 | 2.0 | 40.0 | 60.0 | 8.1 | 93.6 | 6.4 | 7.3 | 100 | 0.0 | 19.3 | 6 | 6 | ||
| fast-pfor | 748 | 95.3 | 49.5 | 11.9 | 24.6 | 14.0 | 35.4 | 4.8 | 2.0 | 86.7 | 13.3 | 4.5 | 79.3 | 20.7 | 3.7 | 87.4 | 12.6 | 20.1 | 6 | 14 | ||
| fileupload | 192 | 100 | 79.7 | 8.9 | 1.0 | 10.4 | 7.8 | 3.6 | 7.3 | 28.6 | 71.4 | 1.6 | 87.5 | 12.5 | 0.0 | 0.0 | 0.0 | 63.4 | 2 | 3 | ||
| graph | 541 | 99.6 | 41.6 | 24.4 | 23.7 | 10.4 | 53.0 | 0.4 | 1.1 | 100 | 0.0 | 1.7 | 100 | 0.0 | 2.2 | 100 | 0.0 | 11.0 | 0 | 1 | ||
| jansi | 409 | 99.8 | 74.6 | 8.1 | 11.5 | 5.9 | 23.5 | 0.2 | 1.7 | 100 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0 | 1 | ||
| pool | 682 | 100 | 75.4 | 5.6 | 2.8 | 16.3 | 21.7 | 1.6 | 1.0 | 100 | 0.0 | 0.3 | 100.0 | 0.0 | 0.0 | 0.0 | 0.0 | 6.6 | 4 | 2 | ||
| validator | 646 | 99.2 | 38.1 | 30.5 | 11.1 | 20.3 | 41.6 | 4.0 | 5.7 | 70.3 | 29.7 | 5.3 | 83.7 | 16.3 | 5.3 | 99.3 | 0.7 | 11.7 | 1 | 15 | ||
4,654 | 99.1 | 43.8 | 25.8 | 16.5 | 13.9 | 38.4 | 3.4 | 3.9 | 67.0 | 33.0 | 7.0 | 69.0 | 11.0 | 3.4 | 66.2 | 3.8 | 15.3 | 30 | 80 | ||
@@
Regardless of GraalVM’s validation for functional equivalence, AlphaTrans translates and executes the test fragments on the recomposed translated project. Super column Test Translation shows the results of test translation and execution. Column TNEF indicates the ratio of AMFs where execution of translated tests never reached them. Columns ATP through ATF show the number of AMFs that AlphaTrans executed using translated test fragments, categorized per the test execution results. On average, for of AMFs, all the test fragments that covered them were marked as pass (ATP). For and , at least one (OTF) or more than one test (MTF) failed. For of them, all the test fragments failed.
@@
For cases with test failure, columns RE and AF show the breakdown of whether test failure was due to assertion failure or runtime error. From the breakdown of results, we can observe that most translated test executions terminated with a runtime error due to translation bugs and never reached the assert statement. Our manual investigation confirms that a high rate of runtime errors is due to a relatively small number of fragments with translation bugs. Although test decomposition helps with test translation coupling effect (§2), there is still a high degree of runtime errors due to the long call chains in these projects (the average number of methods executed per test in the original and decomposed test suites are and ). As a result, the overall pass rate, i.e., the percentage of recomposed test fragments for the translated projects that pass (column TPR), is low.
@@
Finally, we calculated the number of AMFs that GraalVM could not execute (numbers under GE column) but translated test fragments exercised (column M1). All indicates the number of AMFs with all passing tests, including test fragment with assert statement, indicating the validation of functional equivalence. Some corresponds to the number of AMFs with at least one passing test, which indicates runtime validation. Overall, test translation can validate the functional correctness and runtime behavior of and fragments that GraalVM could not exercise.
@@ 7.3. RQ2: Translation Bugs and Fixes
This research question presents the manual effort for fixing translation bugs in a subset of studied subjects, namely Commons-FileUpload, Commons-CLI, Commons-CSV, and Commons-Validator. We also discuss some of the translation bugs and fixes for them to better illustrate the challenges in code translation.
@@ 7.3.1. Human Study
For Commons-FileUpload, achieving green tests took roughly hours and required line additions and line deletions from partial translation generated by AlphaTrans. For Commons-CLI, the manual fix took roughly hours to fully achieve all passing tests, making and line additions and deletions, respectively. For Commons-CSV, the project was very dense, with a lot of method calls, making it harder to manually fix bugs. Nevertheless, a developer achieved all green tests in hours with and line additions and deletions, respectively. Finally, for Commons-Validator, the developer spent hours to fix translation bugs, with and line additions and deletions, respectively. One of the major feedback from developers was that test decomposition greatly helped locate and fix translation bugs: in case of a test failure, developers only need to investigate the last call statement in the failed test fragment instead of looking at the stack trace and other prior calls (more quantified details about the impact of test decomposition on validation in §7.4).

### v3

@@ 7. Empirical Evaluation
Translation Bugs and Fixes. How much effort do developers spend completing the partial translations by AlphaTrans? What is the nature of translation bugs?
@@
Impact of Test Decomposition. What is the impact of test decomposition on validation results?
@@
RQ5:
Ablation Study. To what extent do program transformation, choice of LLM, and program decomposition impact the performance of AlphaTrans?
@@ 7.1. Experiment Setup
We followed three steps for selecting subjects: 1- Mining: We mined GitHub and retrieved a list of repositories that use Java as the primary language, are self-contained (include build files, etc.), and have more than stars with at least one commit pushed within the last months. 2- Filtering: We filtered out projects based on the number of call edges in their call graphs: removed those with less than call edges to ensure the subject projects are big enough to challenge AlphaTrans. We also removed those with more than call edges to reduce the computation and manual effort for further steps. Per GraalVM requirements, we only selected projects we could successfully build (compile and achieve green tests) using Java at . 3- Reduction: AlphaTrans currently supports the following Java APIs: core Java API (java.util, java.text, java.lang, java.io, java.nio, java.net, java.time, and java.math) and third-party libraries (org.opentest4j, org.slf4j.Logger, and org.junit). We automatically removed all other third-party library dependencies and their usage in the source code in the selected projects. We chose a project if at least of its total methods were preserved after such process. Table 1 shows the list of ten projects used in the evaluation of AlphaTrans and details about their size (classes, methods, tests, and fragments).
AlphaTrans uses CodeQL (GitHub, 2024) and tree-sitter (Tree-Sitter, 2024) for static analysis. For running tests, validating translation, and computing coverage, AlphaTrans uses GraalVM (Oracle, 2024), JUnit 4 and 5 (Team, 2024b), Pytest 8.2.1 (pytest dev, 2024), JaCoCo (Team, 2024a), and Python’s coverage (Batchelder, 2024). AlphaTrans works with API- and open-access LLMs. We considered the following criteria for selecting the LLM: (1) for better reproducibility, we prioritized open-access models; (2) due to computing constraints, we wanted an LLM with moderate size ( but parameters); (3) the model should perform reasonably well in code-related tasks; and (4) the model should have fast inference time due to the huge number of prompts. Per the mentioned criteria, we selected DeepSeek-Coder-33b-Instruct for main experiments (RQ1–RQ4). We also used GPT-4o, one of the best-performing commercial models, in RQ5 (§7.6) to demonstrate the impact of stronger models on improving the performance of AlphaTrans. We prompted models under the temperature setting for reproducibility and used their default setting for other parameters. For the base prompting (Algorithm 2), we set the minimum and maximum values of the reprompting budget to and . For the feedback prompting, we set the reprompting budget to , i.e., AlphaTrans attempts to fix issues with feedback only once.
@@ 7.2.1. Effectiveness in type resolution and skeleton validation:
| # Fragments | ||||||||||
| Fields | ||||||||||
| Subjects | # Classes | # Methods | # JUnit Tests | Method Coverage (%) | ATR (%) | SV (%) | Application | Test | Application Methods | Test Methods |
| cli (Foundation, 2024a) | 58 | 664 | 437 | 94.14 | 96.60 | 100 | 104 | 57 | 273 | 2180 |
| codec (Foundation, 2024b) | 156 | 1780 | 992 | 91.03 | 96.01 | 100 | 425 | 140 | 680 | 2849 |
| csv (Foundation, 2024c) | 41 | 694 | 309 | 90.64 | 92.34 | 100 | 146 | 35 | 235 | 1272 |
| exec (Foundation, 2024d) | 56 | 407 | 70 | 54.84 | 78.90 | 100 | 104 | 27 | 248 | 327 |
| fast-pfor (Lemire, 2024) | 82 | 971 | 82 | 54.55 | 87.40 | 100 | 127 | 14 | 748 | 302 |
| fileupload (Foundation, 2024e) | 49 | 381 | 39 | 13.02 | 98.31 | 100 | 121 | 39 | 192 | 194 |
| graph (Foundation, 2024f) | 118 | 879 | 146 | 58.78 | 97.13 | 100 | 216 | 29 | 541 | 975 |
| jansi (Source, 2024) | 48 | 474 | 107 | 23.47 | 84.83 | 100 | 378 | 0 | 409 | 123 |
| pool (Foundation, 2024g) | 98 | 1097 | 73 | 22.29 | 91.60 | 100 | 203 | 91 | 682 | 649 |
| validator (Foundation, 2024h) | 130 | 1228 | 464 | 63.31 | 95.51 | 100 | 421 | 209 | 646 | 1463 |
| Total | 836 | 8575 | 2719 | 56.57 | 91.99 | 100 | 2245 | 641 | 4654 | 10334 |
@@ 7.2.2. Effectiveness in compositional translation and validation:
Table 2 shows the compositional translation and validation results. The AMF column indicates the total number of application method fragments. The numbers in subsequent columns demonstrate the effectiveness of AlphaTrans in the translation and validation of AMFs only444Due to space limit, please refer to our artifact (Lab, 2025) for details of all the translation and validation of other fragments shown in Table 1.. The Syntax Check column indicates the percentage of AMFs that pass syntactic validation. Column SNEF shows the percentage of AMFs not covered by JUnit tests. AlphaTrans successfully generates syntactically correct code ( of AMFs and of all fragments—field, application, and test fragments—across ten subjects). We also observe that of AMFs are not covered during the execution of any test, i.e., we cannot go beyond syntactic check and validate their runtime behavior or functional equivalence.
@@
For of AMFs that can be covered by source project tests, AlphaTrans attempts to validate the functional equivalence using GraalVM. The super column GraalVM shows the percentage of AMFs that GraalVM executes and successfully validates (GS), executes but there is a test assertion failure (GF), and cannot execute due to its limitation (GE) mentioned in §6.1. (min= and max=), (min= and max=), and (min= and max=) of AMFs resulted in Graal Success, Graal Fail, and Graal Error, respectively. Note that these numbers add up to of AMFs that were covered by source program tests. With respect to only covered methods, GraalVM Success will be . Furthermore, our analysis shows that a high portion of methods that are not covered by tests are either abstract methods or getter/setter methods. If their translations are syntactically correct, they are also likely functionally equivalent, which can ramp up the success rate (§7.5). Spearman Rank Order Correlation (Spearman, 1961) indicates a strong positive correlation between the method coverage (Table 1) and GS numbers (), confirming that with better method coverage, the validated AMFs are very likely to be higher.
| GraalVM | Test Translation | |||||||||||||||||||
| OTF (%) | MTF (%) | ATF (%) | M1 | |||||||||||||||||
| Subjects | AMF | Syntax Check (%) | SNEF (%) | GS (%) | GF (%) | GE (%) | TNEF (%) | ATP (%) | O | RE | AF | O | RE | AF | O | RE | AF | TPR (%) | All | Some |
| cli | 273 | 100 | 5.86 | 70.70 | 11.72 | 11.72 | 35.90 | 8.42 | 10.62 | 51.72 | 48.28 | 32.23 | 91.07 | 8.93 | 6.96 | 100 | 0 | 10.08 | 0 | 16 |
| codec | 680 | 98.53 | 8.97 | 38.38 | 32.50 | 20.15 | 65.59 | 4.12 | 4.41 | 60.00 | 40.00 | 12.79 | 55.11 | 44.89 | 4.12 | 75.21 | 24.79 | 9.43 | 11 | 27 |
| csv | 235 | 98.72 | 9.36 | 38.72 | 26.81 | 25.11 | 74.47 | 0 | 13.62 | 96.88 | 3.13 | 0 | 0 | 0 | 2.55 | 100 | 0 | 0 | 0 | 3 |
| exec | 248 | 100 | 45.16 | 33.47 | 2.02 | 19.35 | 34.27 | 4.44 | 2.02 | 40.00 | 60.00 | 7.26 | 93.36 | 6.64 | 6.85 | 100 | 0 | 19.29 | 6 | 9 |
| fast-pfor | 748 | 95.32 | 45.45 | 12.03 | 24.20 | 18.32 | 41.71 | 4.28 | 1.74 | 84.62 | 15.38 | 3.74 | 79.23 | 20.77 | 3.07 | 85.88 | 14.12 | 20.08 | 6 | 25 |
| fileupload | 192 | 100 | 86.98 | 8.85 | 1.04 | 3.13 | 1.56 | 3.65 | 6.77 | 30.77 | 69.23 | 1.04 | 91.67 | 8.33 | 0 | 0 | 0 | 63.44 | 2 | 3 |
| graph | 541 | 99.63 | 41.22 | 24.77 | 22.92 | 11.09 | 57.12 | 0 | 0.92 | 100 | 0 | 0.18 | 100 | 0 | 0.55 | 100 | 0 | 11.04 | 0 | 1 |
| jansi | 409 | 99.76 | 76.53 | 8.07 | 11.49 | 3.91 | 22.25 | 0.24 | 0.98 | 100 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 1.07 | 0 | 1 |
| pool | 682 | 100 | 77.71 | 6.01 | 1.32 | 14.96 | 19.35 | 1.61 | 1.03 | 100 | 0 | 0.29 | 100 | 0 | 0 | 0 | 0 | 6.62 | 4 | 2 |
| validator | 646 | 99.23 | 36.69 | 30.50 | 11.15 | 21.67 | 46.44 | 3.25 | 5.42 | 71.43 | 28.57 | 4.18 | 84.50 | 15.50 | 4.02 | 99.12 | 0.88 | 11.70 | 1 | 31 |
| Total | 4654 | 98.80 | 43.43 | 24.50 | 16.23 | 15.84 | 41.92 | 2.88 | 3.72 | 70.52 | 29.48 | 5.44 | 82.77 | 17.23 | 2.62 | 92.86 | 7.14 | 9.76 | 30 | 118 |
@@
Regardless of GraalVM’s validation for functional equivalence, AlphaTrans translates and executes the test fragments on the recomposed translated project. Super column Test Translation shows the results of test translation and execution. Column TNEF indicates the ratio of AMFs where execution of translated tests never reached them. Columns ATP through ATF show the number of AMFs that AlphaTrans executed using translated test fragments, categorized per the test execution results. For of AMFs, all the test fragments that covered them were marked as pass (ATP). For and , at least one (OTF) or more than one test (MTF) failed. All the test fragments failed for of them. Note that these numbers (TNEF, ATP, OTF, MTF, and ATF) add up to of AMFs covered by source program tests.
@@
For cases with test failure, columns RE and AF show the breakdown of whether test failure was due to assertion failure or runtime error. We can observe that most test executions terminated with a runtime error due to translation bugs and never reached the assert statement. Our manual investigation confirms that a high rate of runtime errors is due to a relatively small number of fragments with translation bugs. Although test decomposition helps with test translation coupling effect (§2), there is still a high degree of runtime errors due to the long call chains in these projects (the average number of methods executed per test in the original and decomposed test suites are and ). As a result, the overall pass rate, i.e., the percentage of recomposed test fragments for the translated projects that pass (column TPR), is low. These results also confirm the necessity of in-isolation testing through language interoperability by AlphaTrans.
@@
We also calculated the number of AMFs that GraalVM could not execute (numbers under GE column) but translated test fragments exercised (column M1). All indicates the number of AMFs with all passing tests, including test fragment with assert statement, indicating the validation of functional equivalence. Some corresponds to the number of AMFs with at least one passing test, which indicates runtime validation. Overall, test translation can validate the functional correctness and runtime behavior of and fragments that GraalVM could not exercise.
Finally, we analyzed translations using PyLint (pyl, 2024), which scores Python files on a scale of to based on how Pythonic the code is. All AlphaTrans translations achieved scores of , mainly because (1) LLMs inherently generate idiomatic code and (2) AlphaTrans uses Black (bla, 2024) for formatting the translations extracted from the LLM response. These results confirm that the translations are all Pythonic, i.e., they adhere to coding standards and best practices.
@@ 7.3. RQ2: Translation Bugs and Fixes
We investigated the manual effort for fixing translation bugs in a subset of studied subjects, namely Commons-FileUpload, Commons-CLI, Commons-CSV, and Commons-Validator. We also discuss some of the translation bugs and fixes for them to better illustrate the challenges in code translation.
@@ 7.3.1. Human Study
For Commons-FileUpload, achieving green tests took hours and required and line additions and deletions from partial translations. For Commons-CLI, the manual fix took hours, making and line additions and deletions, respectively. The project was very dense for Commons-CSV, with many method calls, making it harder to manually fix bugs. Nevertheless, a developer achieved all green tests in hours with and line additions and deletions, respectively. Finally, for Commons-Validator, the developer spent hours to fix translation bugs, with and line additions and deletions, respectively. One of the major feedback from developers was that test decomposition greatly helped locate and fix translation bugs: in case of a test failure, developers only need to investigate the last call statement in the failed test fragment instead of looking at the stack trace and other prior calls.